. Find the area under the standard normal distribution which lies between Z = -0.45 and Z = 0.75?
1
Expert's answer
2020-05-24T15:10:54-0400

Solution: solve this with Laplace function "\Phi" and using a table of Laplace function values.

"\int_{-0.45}^{0.75} \frac{1}{\sqrt{2\pi}}e^{\frac{-x^2}{2}} = \Phi(0.75)-\Phi(-0.45)\approx0.2734+0.1736=0.457"


Answer: 0.457